a. Compute the future value of $2,000 compounded annually for 20 years at 6 percent. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Future value $ 4,207.14 b. Compute the future value of $2,000 compounded annually for 15 years at 9 percent. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Future value $ c. Compute the future value of $2,000 compounded annually for 25 years at 6 percent. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

a

Future value = present value*(1+ rate)^time
Future value = 2000*(1+0.06)^20
Future value = 6414.27

b

Future value = present value*(1+ rate)^time
Future value = 2000*(1+0.09)^15
Future value = 7284.96

c

Future value = present value*(1+ rate)^time
Future value = 2000*(1+0.06)^25
Future value = 8583.74